Lactic acid, the substance that causes sour milk to taste as it does, has a formula weight of 74. It consists of 48.63% C and 8.18% H, with the rest being oxygen. Calculate the empirical and molecular formulas of lactic acid, arranging the atomic symbols in alphabetical order.
1
Expert's answer
2019-06-19T14:14:48-0400

Since C= 48.63 %, thus no. of carbon atoms can be calculated as 48.63/12= 4.05

Similarly, for H = 8.18/1 = 8.18

For Oxygen (100-48.63-8.18)/16 = 2.69


Thus, the formula becomes C4.05H8.18O2.69

Thus, dividing each entity by 2.69 The formula becomes: C1.5H3.0O1 (Empirical Formula)

Multiplying above formula by 2 to get whole number ratio gives: C3H6O2 (molecular formula)